OpenStudy (anonymous):

Systems of equations with the same slope but diffirrent y interceps have no solution Always right

OpenStudy (anonymous):

SOMETIMES they do, not always.

OpenStudy (amistre64):

a solution is determined by where the equations equal each other spose you take 2 equations with the same slope and different y intercepts y = ax + b y = ax + c and equate them; y=y ax + b = ax + c we can subtract ax from each side giving us b = c ; but this cannot be true since they have different y intercepts, therefore b does nto equal c